OpenStudy (anonymous):

show that if n is even then \(2^n-1 \) divides 3

hartnn (hartnn):

let n =2k

hartnn (hartnn):

\(\Large 4^k-1 = (4-1)(...)\) done :)

OpenStudy (anonymous):

haha nice

hartnn (hartnn):

k belongs to integers, right ?

OpenStudy (anonymous):

yeah lol 100%

OpenStudy (anonymous):

mark why your cat is SO CUTE @Marki

OpenStudy (anonymous):

ikr :)

OpenStudy (michele_laino):

anothe demostration can be made using the induction principle, namely, if I set n=2k, where k is a positive integer greater than zero, then I have: \[k=1\rightarrow 2^{2k}-1=4-1=3\] so our statement is true, \[k=2\rightarrow 2^{2k}-1=16-1=15\] so our statement is true again. Now I suppose that our statement is true for a generic k, and I write that statement for k+1, namely: \[2^{2(k+1)}-1=2^{2k}*4-4+4-1=(2^{2k}-1)*4+3\] so our statement, by the induction principle is true again, and it is true for all k, or for all even n